David's Astronomy Pages
Notes - Session 1305 (2025-03-23)

 
Bullet Session Aims & Highlights
 - Observing Result
 - Night Summary Plot
 - Session Event Log
Bullet Operational Issues
  - Critical Issues (0),  Major Issues (1),  Minor Issues (1),  Small Defects (1),  Continuous Improvement (2)
Bullet Images from 2025-03-23 >>         
Bullet Geomagnetic Station - Site Prep (2025-03-23 to 2025-03-25)
Bullet Geomagnetic Station - Assembly Design (2025-03-26 to 2025-03-27)
   

Session Aims & Highlights (2025-03-23)

Main aims

  1. Targets.  Acquire images of a selection of variable stars, nearby stars, comets & deep sky targets as allowed by sky conditions.

 Equipment & Software

Highlights

Notes

Summary Plots & Logs

Observing Plan
Image
  
Observing Result
Image
   
Image
 
Dome & Scope Slewing Performance
Image
  
Slew/Centering Performance
Image
  
Guiding Performance
Image
  
Sky Conditions (Locate Frames)
Image 
  
Night Sky Summary Plot
Top axis: Sky Brightness at Zenith (in ADU/s)
Lefthand axis: Local Time (hh LT). Righthand axis: Sun Altitude (degs)
Image   
  
Actual Weather vs Pre-Session Weather Forecast
Image
Image
  
Session Event Log
Time     Event Detail
22:56:18 Session Monitoring AutoStart monitoring for Live Session opportunity between 22:56 & 03:53
22:56:22 Session AutoStarting Session autostarting (22:56)
22:56:45   CCDSoft Restarting CCDSoft being restarted (to set AutoSave No.)
22:57:31   Camera1 Connected SBIG Camera connected (Set point -25°C)
22:57:33 Session Created Session Created (Live, 2025-03-23 S01305, ImageSaveNum: 1305001)
22:57:40   Scope Switched On Telescope Power has been switched on via UPB Powerbox.
22:59:23   Services Started Observatory Services started
22:59:31 Observatory (Auto) Observatory placed in Fully-Automated Mode
22:59:35 Session Pending Session pending (2025-03-23)
22:59:37 Session Initiating Session initiating (2025-03-23)
22:59:46   Camera1 Connected SBIG Camera connected (Set point -20°C)
22:59:54   Plan Requested Observing Plan requested from AstroPlan (1.43)
23:00:35   Plan Loaded Observing Plan loaded to queue (Plan ID: 1089)
23:01:16   Telescope Connected Telescope connected (TheSky6)
23:01:40 Session Equilibration Session ready to Open Dome
23:02:25   Dome Opened Dome opened (Opening time 46s, Zigbee 41s)
23:02:55 Session Running Session running
23:02:58   Queue Started Observing Queue started (16 targets selected)
23:03:00     Target Started (NrZen) Target started (Focus Field 11, HIP 53798)
23:06:48       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
23:11:07       Focusing Completed Foc1 AutoFocus Completed (Profile No 1, wide)
23:13:31       Focusing Completed Foc1 AutoFocus Completed (Profile No 1)
23:13:34       Focusing Started-Foc2 Foc2 Focusing Started (Secondary Scope, using ShCap)
23:15:39       Focusing Completed Foc2 AutoFocus Completed (Profile No 2, wide)
23:17:33       Focusing Completed Foc2 AutoFocus Completed (Profile No 2)
23:17:52     Target Completed (NrZen) Target completed (Focus Field 11, HIP 53798)
23:17:56     Target Started (1/16) Target started (1/16, UGC 5695 w/SN2025cfc)
23:23:21   Pulsar Dome Restarting Dome Controller
23:23:58       SoftSuspend Called Soft Suspend called due to Deteriorating Conditions (low battery)
23:24:01   Pulsar Dome Dome Controller has been restarted
23:24:08       Focusing Skipped Foc1 focusing skipped - unable to find a star (TCF-S)
23:24:14   Pulsar Dome Dome Service has been restarted
23:24:18       SoftSuspend Cancelled Soft Suspend is cancelled due to improved conditions
23:24:25       Dome Unparked Dome unparked
23:25:01   Pulsar Dome Recovering Dome
23:25:13   Pulsar Dome Restarting Dome Controller
23:25:48   Pulsar Dome Dome Controller has been restarted
23:25:53       SoftSuspend Called Soft Suspend called due to Deteriorating Conditions (low battery)
23:26:00   Pulsar Dome Dome Service has been restarted
23:26:03       SoftSuspend Cancelled Soft Suspend is cancelled due to improved conditions
23:26:37   Telescope Connected Telescope connected (TheSky6)
23:27:36 User Intervention User Intervention to abort job queue & switch back to Manual Mode
23:27:39     Target Completed Target partially completed (1/16, UGC 5695 w/SN2025cfc)
23:27:41   Queue Aborted Job Queue aborted
23:28:28   Services Started Observatory Services started
23:29:38   Services Stopped Observatory Services stopped
23:29:44 Session Closed Session closed by User
23:29:47 Program Closed Program closed by User
23:30:56 Program Opened Program Opened (AstroMain 3.73.14)
23:30:58   CCDSoft Restarting CCDSoft being restarted (to set AutoSave No.)
23:31:11 Session Resumed Session Resumed (Live, 2025-03-23 S01305, ImageSaveNum: 1305007)
23:31:17   Obs.Overseer Started Obs.Overseer started
23:33:00   Obs.Manager Started Obs.Manager started
23:33:09   Services Started Observatory Services started
23:35:15 Session Monitoring AutoStart monitoring for Live Session opportunity between 23:35 & 03:53
23:35:17 Observatory (AutoStart) Observatory placed in Auto-Start Mode
23:35:19 Session AutoStarting Session autostarting (23:35)
23:35:37   Services Started Observatory Services started
23:35:45 Observatory (Auto) Observatory placed in Fully-Automated Mode
23:35:48 Session Pending Session pending (2025-03-23)
23:35:50 Session Initiating Session initiating (2025-03-23)
23:35:57   Camera1 Connected SBIG Camera connected (Set point -20°C)
23:36:05   Plan Loaded Observing Plan loaded to queue (Plan ID: 1089)
23:36:15   Telescope Connected Telescope connected (TheSky6)
23:36:37 Session Running Session running
23:36:40   Queue Started Observing Queue started (15 targets selected)
23:38:42     Target Started (2/16) Target started (2/16, NGC 2744 w/SN2025oq)
23:42:14       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
23:45:10       Focusing Completed Foc1 AutoFocus Completed (Profile No 5)
00:01:23     Target Completed Target completed (2/16, NGC 2744 w/SN2025oq)
00:01:27     Target Started (3/16) Target started (3/16, 29P/Schwassmann-Wachmann)
00:04:58       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
00:07:46       Focusing Completed Foc1 AutoFocus Completed (Profile No 6)
00:23:20     Target Completed Target completed (3/16, 29P/Schwassmann-Wachmann)
00:23:24     Target Started (4/16) Target started (4/16, NGC 3277 w/SN2025coe)
00:26:56       Focusing Skipped Foc1 focusing skipped - star is lost (TCF-S)
00:43:03     Target Completed Target completed (4/16, NGC 3277 w/SN2025coe)
00:44:36     Target Started (5/16) Target started (5/16, AT2025euu)
00:48:12       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
00:50:53       Focusing Completed Foc1 AutoFocus Completed (Profile No 7)
01:07:27     Target Completed Target completed (5/16, AT2025euu)
01:07:31     Target Started (6/16) Target started (6/16, M63)
01:11:37       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
01:14:07       Focusing Completed Foc1 AutoFocus Completed (Profile No 8)
01:30:17     Target Completed Target completed (6/16, M63)
01:30:22     Target Started (7/16) Target started (7/16, T CrB)
01:33:57       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
01:36:37       Focusing Completed Foc1 AutoFocus Completed (Profile No 9)
01:41:46     Target Completed Target completed (7/16, T CrB)
01:41:51     Target Started (8/16) Target started (8/16, NGC 6384 w/SN2024pxl)
01:45:50       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
01:48:45       Focusing Completed Foc1 AutoFocus Completed (Profile No 10)
02:04:59     Target Completed Target completed (8/16, NGC 6384 w/SN2024pxl)
02:05:03     Target Started (9/16) Target started (9/16, AT2022uam (CV))
02:08:18       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
02:10:50       Focusing Completed Foc1 AutoFocus Completed (Profile No 11)
02:16:49     Target Completed Target completed (9/16, AT2022uam (CV))
02:16:53     Target Started (10/16) Target started (10/16, AT2025esv)
02:19:51       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
02:22:33       Focusing Completed Foc1 AutoFocus Completed (Profile No 12)
02:38:49     Target Completed Target completed (10/16, AT2025esv)
02:38:53     Target Started (11/16) Target started (11/16, AT2025oj (CV))
02:42:03       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
02:44:56       Focusing Completed Foc1 AutoFocus Completed (Profile No 13)
03:00:55     Target Completed Target completed (11/16, AT2025oj (CV))
03:00:59     Target Started (12/16) Target started (12/16, GCVS BL Lac)
03:04:12       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
03:06:42       Focusing Completed Foc1 AutoFocus Completed (Profile No 14)
03:11:13     Target Completed Target completed (12/16, GCVS BL Lac)
03:11:17     Target Started (13/16) Target started (13/16, C/2023 A3 (Tsuchinshan-ATLAS))
03:15:20       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
03:18:24       Focusing Completed Foc1 AutoFocus Completed (Profile No 15)
03:25:13     Target Completed Target completed (13/16, C/2023 A3 (Tsuchinshan-ATLAS))
03:25:18     Target Started (14/16) Target started (14/16, 2MFGC 13744 w/SN2024afyu)
03:29:29       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
03:32:01       Focusing Completed Foc1 AutoFocus Completed (Profile No 16)
03:48:38     Target Completed Target completed (14/16, 2MFGC 13744 w/SN2024afyu)
03:48:42     Target Started (15/16) Target started (15/16, UGC 10913 w/AT2025hy)
03:52:16       Focusing Started-Foc1 Foc1 Focusing Started (TCF-S)
03:54:35       Focusing Completed Foc1 AutoFocus Completed (Profile No 17)
04:11:05     Target Completed Target completed (15/16, UGC 10913 w/AT2025hy)
04:11:09     Target Started (16/16) Target started (16/16, AT2025esx)
04:30:10     Target Completed Target completed (16/16, AT2025esx)
04:30:13   Queue Completed Job Queue completed
04:30:16 Session Closing Session closing
04:31:11   Dome Closed Dome closed (Closing time 50s, Zigbee 44s)
04:31:51   Dome Parked Dome parked (parking time 33s), Az: 90.0 deg
04:32:30   Telescope Parked Telescope parked (parking time 32s total)
04:33:00   Telescope State Handbox reads 'Scope parked Turn scope off.'
04:33:22   Telescope Switched Off Telescope Power has been switched off via UPB Switch.
04:34:48   Services Stopped Night Services stopped
04:34:51 Session Finishing Session Finishing started (Create Fits Summary, Transfer Files)
04:35:13     Dome (Find Park) Find Park started (Search Az 88.0 to 92.0°, Step 0.2°, Narrow)
04:35:57     Dome Find Park found (Best Park Az 90.6°)
04:36:03     Dome Dome has been synced at park position (Az 90.0°, Adjustment: -0.6°)
04:36:05   Dome Parked Dome parked
04:38:16   Dome Charging Dome is parked and re-charging Ok (19mA)
04:38:19 Session Finished Session Finished
 
Session Alerts & Alarms
Time     Type       Name Detail
22:59:35 Yellow Alert Disk Space Free Space on C:\ Drive (36.20 GB) is less than 50 GB
23:18:40 Yellow Alert Dome Discrepency Difference between Dome Az and Telescope Az is > 8 deg (UGC 5695)
23:35:46 Yellow Alert Disk Space Free Space on C:\ Drive (36.85 GB) is less than 50 GB
02:39:30 Yellow Alert Dome Dome position has been static for more than 5 mins
04:35:09 Yellow Alert Disk Space Free Space on C:\ Drive (37.02 GB) is less than 50 GB
04:38:17 Green Clear Dome Charging Dome is parked and re-charging Ok (19mA)
 

Back to Top


Operational Issues (2025-03-23, S1305)

[ Prev | Next ]

Critical Issues

Major Issues

Minor Issues

Small Defects

Continuous Improvement

[ Prev | Next ]


Fig 1.  Timeline associated with Slaved Dome Anomaly at start of session

P   23:01:40.475 Dome Commands: OpenShutter: (slew started)

P   23:01:40.475 OpenShutter Request sent

P   23:01:40.475 ShutterStatus 2
P   23:01:40.475 Dome Statuses: Get ShutterStatus: shutterOpening

P   23:02:23.264 ShutterStatus 0

CL  23:03:09.16 | Focus Field 11                         Call    Scope.SlewToRaDec   Ok      | Calling TheSky6.Telescope.SlewToRaDec
C1  23:03:09.553 Get Slaved: False (done)
C1  23:03:09.553 Set Slaved: Setting Slaved to True(Enable slaving for ASCOM.Pulsar_Observatories_Dome.Dome) (done)
CL  23:03:09.55 | Focus Field 11                         Set     Dome.Slaving(On)    Info    | Set Dome Slaving ON
C1  23:03:10.425 Get Slaved: True (done)
DH 23:03:10.969 Telescope Commands: SlewToCoordinatesAsync - RA = 11:01:29.3, Dec 50:54:03: (slew started)

DH 23:18:05.789 Telescope Commands: SlewToCoordinatesAsync - RA = 10:30:59.5, Dec 13:01:49: (slew started)

DH 23:18:06.525 Telescope Statuses: Get Slewing: True
DH 23:18:32.090 Telescope Statuses: PollScopeTask - Cancelling slew in progress state
DH 23:18:33.530 Telescope Statuses: Get Slewing: False

C1 23:23:20.473 Get Slaved: True (done)
DH 23:21:52.162 Dome Commands: OpenShutter: (slew started)
P    23:21:52.162 OpenShutter Request sent

P    23:24:22.611 SlewToAzimuth Start 200

P    23:24:22.617 Slewing Get False (Meaning True)

P    23:24:23.552 Azimuth Get Azimuth = 90
P    23:24:23.552 AtPark Get AtPark = True

P    23:24:24.476 Azimuth Get Azimuth = 91.7
P    23:24:24.477 AtPark Get AtPark = False

P    23:25:06.155 Azimuth Get Azimuth = 185.2
 - Dome Restarted

P    23:25:48.884 CommandString RX 200.0
P    23:25:50.465 Azimuth Get Azimuth = 200

C1    23:38:50.471 Set Slaved: Setting Slaved to True(Enable slaving for ASCOM.Pulsar_Observatories_Dome.Dome) (done)

 

[ Prev | Next ]

Back to Top


Geomagnetic Station - Site Prep (2025-03-23 to 2025-03-25)

A Geomagnetic Monitoring Station is being added to the Observatory in Spring 2025 for monitoring changes in the earth's magnetic field and to supplement and build upon the Observatory's existing Aurora monitoring that use a Black/White AllSky Camera and new NorthCam Colour Camera.

The following notes describe the Site Preparation Phase

Other Notes  :
| Review | SAM III | Site Prep | Assembly Design | Parts Assembly |  Testing  | Final Assembly | Installation | Commissioning | Remedial Work |

Hole Digging (2025-03-23)

Hole that will contain the 3-axis geomagnetometer sensor fixture has been commenced on 2025-03-21 and reached 45 cm depth.  Pictures below show status of hole on 2025-03-23 after reaching 68cm.

Hole that will eventually contain
3-axis geomagnetometer sensor fixture

Hole has reached 68cm deep at the time of the photo
another 52cm to go in red subsoil / drift deposit
    Hole location relative to Observatory
where the SAM III Monitor will reside
Offset Distance 10m
Cables will eventually be run below the lawn.
Image   Image
      

Hole Digging (2025-03-24)

Hole that will contain the 3-axis geomagnetometer sensor fixture has now reached 100cm depth.  Another 20cm to go

Hole Completion (2025-03-25)

Hole that will contain the 3-axis geomagnetometer sensor fixture has been completed after reaching 120cm depth. 

Completed Hole for the
3-axis geomagnetometer sensor fixture
Hole depth is 120cm
    Detailed image showing lower part of hole
Image   Image
      
Side view of hole
120cm depth
    Side view of hole
120cm depth
.
Image   Image
     

| Review | SAM III | Site Prep | Assembly Design | Parts Assembly |  Testing  | Final Assembly | Installation | Commissioning | Remedial Work |

Back to Top


Geomagnetic Station - Assembly Design (2025-03-26 to 2025-03-27)

A Geomagnetic Monitoring Station is being added to the Observatory in Spring 2025 for monitoring changes in the earth's magnetic field and to supplement and build upon the Observatory's existing Aurora monitoring that use a Black/White AllSky Camera and new NorthCam Colour Camera.

The following notes describe the Assembly Design & Parts.

Other Notes  :
| Review | SAM III | Site Prep | Assembly Design | Parts Assembly | TestingFinal Assembly | Installation | Commissioning | Remedial Work |


Pipe work for the building the fixture that will house the 3-axis geomagnetometer sensors & the temperature sensor were purchased today (2025-03-26) from a local hardware shop.   Items comprised FloPlast 40/43 mm diameter straight pipe, angled junctions, connectors and end stops. Junctions will eventually be joined/sealed using special FloPlast solvent cement.

A mock-up of the fixture showing the rough positions of sensors is shown in the picture below.  Note that the straight pipe has not been cut to size yet, and the top horizontal sensors will be placed at 90 deg to each other in the final build (not the 180 deg positioning used in the rough mock-up).  Cables to the magnetometer sensors aren't shown here,but the Temperature Sensor & its cable are.


Image

Update 2025-03-27 (Sensor Fixture Progress)

Pipe work for the building the buried sensor fixture has been cut to length and the entire assembly loosely fitted together (see pictures below) :


   Image

3 x 15m lengths of CAT6 Cable have been cut from 100m reel previously purchased. 

The required cable length is around 12.5m, so these 15m lengths have an excess which will be trimmed off during the final tieback to the Observatory.

| Review | SAM III | Site Prep | Assembly Design | Parts Assembly | TestingFinal Assembly | Installation | Commissioning | Remedial Work |

 

Back to Top